Tech Staff Meeting Notes August 16, 2004

Current Top Priorities are:BB Upgrade, TZTN


Greg's list

Monster south campus run August 23.

Sunlab. done.

Routine Maintenance. DBB PM's.

BB network upgrade. VHF transceivers failed and have to be redesigned and rebuilt.

Battery charging station. Need a hood. What to do about a body wash?

Still need to finish 4th channel up grade on two stations. done. Need to repair several that are apparently dead. done. Something flakey with the marked tree gain channels.

Few more strongmotion stations left for routine service.


Need to get malfunctioning systems back in order to install.
Need to revisit getting continuous connections to all possible stations.
Parker was going to CVVA wednesday to upgrade but got rerouted to SWET for repair

I-40 Bridge. Are we going to use TDOT's ITS system for comm and if so can we send aztn to the bridge and avoid dsl? Would supposedly supply a connection to the university. How do you get the rest of the way to CERI. Would need to put two pc's at the access point. Will this project ever stop changing? Software development continues.

MKTA AC. Will replace.

FDL. Current scheme is to get Horton involved to provide direction for organizing, inventorying, etc.

Second Tier Priorities

FDL flooding. Knock down that wall.

CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. Filipcic needs to install.

MPHB. Mitch needs to determine current orientation so we can pull it and orient correctly.

BB Test Facility. Need to pour the slabs, install the conduit and electrical enclosure, etc. Now that soil voids are taken care of. Now what to do about the voids between the bottom of the vault and the concrete? Should we rent a tool to "settle" the ground around the vault (not this project does not have a budget).

VBMS. Have an agreement for signatures. Need to finish proposal for installation funds.

SEKY (now TZTN). Need to submit proposal in June for installation by 9/30. Proposal submitted to IRIS. Need to find out what DAS/sensor package we're installing. Mac/Parker will arrange prelim electrical work. Mitch will get more info from Golden on hardware to be installed and see about getting here or Golden folks scheduled to install.

Timing. Greg has an easy way to generate a datastream that goes high when lock is lost. Need a worm module to monitor. BB monitoring is more difficult. How do we automate the status stream or even get it back to Memphis for logging? (Save it to file with Scream?). Note, new version of firmware sends status to both flash and stream. Hence, information will be in gcf2ew logs, but how do we monitor that? Maybe not a huge deal with broadbands since both Jim and Steve regularly monitor. Would still be good to have a record though.

Cowcamps pit is done. A tower is being constructed at the top of the mountain that have permission to get on.

Blohole property now in the hands of a private landowner. Will allow continued operation of the station but no vehicles (i.e. a 1.5 mile hike). John as some leads on an alternative site.


North Carlina station refurbish. Keep node at LRNC. Modify to be the same as a New Madrid node and minimize internet bandwidth required.

Pulling SMA's as convenient. Hardware to be cannibalized. Done.

Calibrate S-13's.

Nagios. Reno uses this to monitor SOH for a variety of systems including nodes and seismic stations. Could be very useful.

Got batteries for GPS stations, need to install. done.

Gural EDU seismometers received. Works fine. Greg is working on power systems. Done?